# Logarithmic delocalization of low temperature 3D Ising and Potts interfaces above a hard floor

### v1

@@ 2.1. Random-cluster, FK–Potts and fuzzy Potts models
    Let $\mathcal{V}_{\mathsf{top}}$ denote the vertices of the open cluster of $\omega$ incident to the $\mathsf{top}$ wired boundary vertices; augment it to $\widehat{\mathcal{V}}_{\mathsf{top}}$ by adding all connected components of $(\mathbb{Z}+\frac{1}{2})^{3}\setminus\mathcal{V}_{\mathsf{top}}$ whose outer veretx boundary is a subset of $\mathcal{V}_{\mathsf{top}}$. Define $\mathcal{V}_{\mathsf{bot}}$, $\widehat{\mathcal{V}}_{\mathsf{bot}}$ analogously.
@@
By the FK–Potts coupling, one has $\widehat{\mathcal{V}}_{\mathsf{top}}\subset\widehat{\mathcal{V}}_{\mathsf{red}}\subset\widehat{\mathcal{V}}_{\mathsf{blue}}^{c}$ and $\widehat{\mathcal{V}}_{\mathsf{bot}}\subset\widehat{\mathcal{V}}_{\mathsf{blue}}(\subset\widehat{\mathcal{V}}_{\mathsf{red}}^{c})$, whence the four interfaces are ordered via the partial order on subsets of the vertices [^2] Associate $\mathcal{I}_{\mathsf{blue}}$ and $\mathcal{I}_{\mathsf{bot}}$ with $\widehat{\mathcal{V}}_{\mathsf{blue}}$ and $\widehat{\mathcal{V}}_{\mathsf{bot}}$, resp., and associate $\mathcal{I}_{\mathsf{red}}$ and $\mathcal{I}_{\mathsf{top}}$ with $\widehat{\mathcal{V}}_{\mathsf{red}}^{c}$ and $\widehat{\mathcal{V}}_{\mathsf{top}}^{c}$, resp. (i.e., in all cases, use the part incident to the bottom face of the boundary conditions) to arrive at this partial order. :
@@
Let $\omega\sim\bar{\pi}_{n}^{\mathsf{fl}}$ or $\omega\sim\bar{\pi}_{n}^{\mathsf{h}}$. The full interface $\mathcal{I}_{\mathsf{full}}$ is the $1$-connected component of plaquettes dual to closed edges of $\omega$ (the plaquettes $f,f^{\prime}$ are $1$-adjacent if they share an edge) that separates the $\mathsf{top}$ and $\mathsf{bot}$ wired boundary components.
@@
The above result is stated in [14, Lem. 9],[15, Lem 7.118] for the limit as $m\to\infty$, but its analog was first established for $\bar{\pi}^{\mathsf{h}}_{\Lambda^{\prime}_{n,m}}(\mathcal{I}_{\mathsf{full}}=I)$ in [14, Lem. 7],[15, Lem 7.104] for all $n,m\geq 1$. The proof further applies to any box $\Lambda=([-\frac{n}{2},\frac{n}{2}]^{2}\times[m_{1},m_{2}])\cap(\mathbb{Z}+\frac{1}{2})^{3}$ for $m_{1}\leq 0$ and $m_{2}\geq 1$, with the caveat that the bound on ${\mathbf{g}}$ in terms of $r(f,I;f^{\prime},I^{\prime})$ (Eq. 2.10) might render Eq. 2.8 useless if $m_{1}$ or $m_{2}$ were taken too close to 00 (as $r$ will be capped at the distance to the boundary).
@@
Let $q\geq 1$ and $n,m\geq 1$. There exist $p_{0}<1$ such that, for every $p>p_{0}$ there is some $a=a_{p}>0$ so the following holds,
@@
The above result was stated for the limit $m\to\infty$, and as explained in Remark 2.4, is further applicable to $\Lambda^{\prime}_{n,m}$. In our context of $p=1-e^{-\beta}$, and for $\beta$ large enough the result above holds true for $a=\beta-C$ where $C>0$ is an absolute constant, provided we have $1\leq k(n)=o(m)$ (see [5, Props. 2.13]). As per Remark 2.4, it also holds in $\Lambda=([-\frac{n}{2},\frac{n}{2}]^{2}\times[m_{1},m_{2}])\cap(\mathbb{Z}+\frac{1}{2})^{3}$ for $m_{1}\leq 0<m_{2}$ such that $k=o(|m_{1}|\wedge m_{2})$ (the effect of which is negligible compared to $e^{-ak}$).
@@ 2.2. A logarithmic upper bound on the top interface
As explained by step (2) in the sketch at the beginning of this section, the lower bound argument requires as an input, a logarithmic upper bound on the maximal height of $\mathcal{I}_{\mathsf{top}}$ in the measure $\bar{\pi}_{n}^{\mathsf{fl}}$. Recall that $\mathcal{L}_{\geq h}$ is the slab $\mathbb{R}^{2}\times[0,\infty)$.
@@
Consequently, by Eq. 2.7, the same holds for $\mathcal{I}_{\mathsf{bot}}$ under $\bar{\pi}_{n}^{\mathsf{fl}}$, as well for $\mathcal{I}_{\mathsf{blue}},\mathcal{I}_{\mathsf{red}}$ under ${\widehat{\upmu}}_{n}^{\mathsf{fl}}$.
While the logarithmic upper bound on $\mathcal{I}_{\mathsf{top}}$ implies one for $\mathcal{I}_{\mathsf{blue}}$, our argument for establishing this bound for $\mathcal{I}_{\mathsf{top}}$ goes through first proving such a result directly on $\mathcal{I}_{\mathsf{blue}}$ (Corollary 2.12 below) and then using the rigidity of the $4$ interfaces (Corollary 2.14 below) to conclude the lemma above.
@@ 2.2.1. Upper bound on the blue interface with a soft floor
By Theorems 2.5 and 2.6, a union bound over $x=(x_{1},x_{2})\in[-\frac{n}{2},\frac{n}{2}]\cap(\mathbb{Z}+\frac{1}{2})^{2}$, as well as on the two events $\{\overline{\operatorname{ht}}_{x}(\mathcal{I}_{\mathsf{full}})\geq h_{0}\}$ and $\{\underline{\operatorname{ht}}_{x}(\mathcal{I}_{\mathsf{full}})\leq-h_{0}\}$ in Theorem 2.5 (which, in our context here of $\bar{\pi}_{n}^{h_{0}}$, address the intersection of $\mathcal{I}_{\mathsf{full}}$ with $\mathcal{L}_{2h_{0}}$ and $\mathcal{L}_{0}$, respectively), gives
@@ 2.2.2. Comparing the soft and hard floor measures
Determining whether $\mathcal{I}_{\mathsf{blue}}\subset\mathcal{L}_{\geq 0}$ must be done with care; a good example to keep in mind in this context is $\sigma$ that colors every $v\in\mathcal{L}_{<0}$ in $\mathsf{blue}$ and every $v\in\mathcal{L}_{>0}$ in $\mathsf{red}$ except for $Q=\{\pm\frac{1}{2}\}^{3}$, where we swap the assignment, coloring $v\in Q\cap\mathcal{L}_{<0}$ in $\mathsf{red}$ and $v\in S\cap\mathcal{L}_{>0}$ in $\mathsf{blue}$. In this example, $[0,1]^{2}\times\{-1\}\subset\mathcal{I}_{\mathsf{blue}}$ (in particular $\mathcal{I}_{\mathsf{blue}}\nsubseteq\mathcal{L}_{\geq 0}$) and similarly $[0,1]^{2}\times\{1\}\subset\mathcal{I}_{\mathsf{red}}$, since $Q\subset\widehat{\mathcal{V}}_{\mathsf{blue}}^{c}\cap\widehat{\mathcal{V}}_{\mathsf{red}}^{c}$. On the other hand, if we were to also set the color of $(\frac{3}{2},\frac{1}{2},\frac{1}{2})$ to $\mathsf{blue}$, it would connect $Q\cap\mathcal{L}_{>0}$ to $\mathcal{V}_{\mathsf{blue}}$, thus leading to $Q\subset\widehat{\mathcal{V}}_{\mathsf{blue}}$ and to $\mathcal{I}_{\mathsf{blue}}=\mathcal{I}_{\mathsf{red}}\subset\mathcal{L}_{\geq 0}$.
[…]
@@
arriving at $\phi_{n}^{\mathsf{fl}}$, a fuzzy Potts measure on $\Lambda_{n}$ with a hard $\mathsf{blue}$ floor at height 00, a single vertex $v_{\mathsf{red}}$ replacing all the $\mathsf{red}$ boundary vertices in $\mathcal{L}_{>\mathsf{h}}$, and that vertex connected to every formerly $\mathsf{blue}$ boundary condition vertex in $\partial\Lambda_{n,n}\cap\mathcal{L}_{>0}\cap\mathcal{L}_{<\mathsf{h}}$ via edges with interaction strength $J$):
@@ 2.2.3. Rigidity of the four interfaces
Let $\omega\sim\bar{\pi}_{n}^{\mathsf{fl}}$, consider $v\in\Lambda_{n,n}$. Let $\mathcal{C}_{v}$ be the (potentially empty) $(\mathbb{Z}+\frac{1}{2})^{3}$-connected component of the vertex $v$ in $\Lambda_{n,n}\setminus(\widehat{\mathcal{V}}_{\mathsf{top}}\cup\widehat{\mathcal{V}}_{\mathsf{bot}})$. Let $\Gamma_{v}$ be its plaquette boundary, that is, $f\in\Gamma_{v}$ if it is dual to the $(\mathbb{Z}+\frac{1}{2})^{3}$-bond $(u,v)$ with $u\in\mathcal{C}_{v}$ and $v\notin\mathcal{C}_{v}$. Then
@@
Fix $\beta$ large enough. For sufficiently large $n$, with $\bar{\pi}_{n}^{\mathsf{fl}}$-probability at least $1-n^{-1+\varepsilon_{\beta}}$ the following holds. For every plaquette $f\in\mathcal{I}_{\mathsf{top}}\setminus\mathcal{I}_{\mathsf{bot}}$ there exists $f^{\prime}\in\mathcal{I}_{\mathsf{bot}}\setminus\mathcal{I}_{\mathsf{top}}$ such that $f,f^{\prime}$ can be connected by a path of adjacent plaquettes in $\Lambda_{n,n}$ of length at most $(8/\beta)\log n$. In particular,
$\bar{\pi}_{n}^{\mathsf{fl}}\left(\max_{x}\overline{\operatorname{ht}}_{x}(\mathcal{I}_{\mathsf{top}})>\max_{x}\overline{\operatorname{ht}}_{x}(\mathcal{I}_{\mathsf{bot}})+(8/\beta)\log n\right)\leq n^{-1+\varepsilon_{\beta}}\,.$
The analogous statement holds for $f\in\mathcal{I}_{\mathsf{bot}}\setminus\mathcal{I}_{\mathsf{top}}$ and $f^{\prime}\in\mathcal{I}_{\mathsf{top}}\setminus\mathcal{I}_{\mathsf{bot}}$, and in particular,
$\bar{\pi}_{n}^{\mathsf{fl}}\left(\min_{x}\underline{\operatorname{ht}}_{x}(\mathcal{I}_{\mathsf{bot}})<\min_{x}\underline{\operatorname{ht}}_{x}(\mathcal{I}_{\mathsf{top}})-(8/\beta)\log n\right)\leq n^{-1+\varepsilon_{\beta}}\,.$
@@
Every $f\in\mathcal{I}_{\mathsf{top}}\setminus\mathcal{I}_{\mathsf{bot}}$ belongs to the boundary $\Gamma_{v}$ of some component $\mathcal{C}_{v}$ as per the definition in Lemma 2.13. Note that said boundary must also contain some plaquette $f^{\prime}\in\mathcal{I}_{\mathsf{bot}}\setminus\mathcal{I}_{\mathsf{top}}$ (some external boundary vertex of $\mathcal{C}_{v}$ must be in $\widehat{\mathcal{V}}_{\mathsf{bot}}$, or else $\mathcal{C}_{v}$ would be part of $\widehat{\mathcal{V}}_{\mathsf{top}}$), whence $|\Gamma_{v}|$ is an upper bound on the length of a path of plaquettes connecting $f,f^{\prime}$. By a union bound over the $O(n^{3})$ vertices in $\Lambda_{n,n}$, we have $\max_{v}|\Gamma_{v}|\leq(8/\beta)\log n$ with probability at least $1-n^{-1+\varepsilon_{\beta}}$, as claimed. The statement for $f\in\mathcal{I}_{\mathsf{bot}}\setminus\mathcal{I}_{\mathsf{top}}$ follows from the exact same reasoning. ∎
@@
Let $(\sigma,\omega)\sim{\mathbf{P}}_{n}^{\mathsf{fl}}$. By Corollary 2.14, combined with the ordering of the interfaces in Eq. 2.7, with probability at least $1-n^{-1+\varepsilon_{\beta}}$ we have that for every $x$,
@@
For any finite graph $(V,E)$ and an edge set $\tilde{E}$, fix $p,q$ and let $\tilde{Z}_{G}^{(\theta)}$ be the partition function of the $(p,q)$-FK model on $G=(V,E\cup\tilde{E})$ modified to have edge probability $\theta$ for $e\in\tilde{E}$. Denote by $\tilde{\pi}_{G}^{(\theta)}$ the probability of an edge configuration $\omega$ in this model. Then
@@
Let $I\subset\mathcal{L}_{\geq 0}$ be a realization for $\mathcal{I}_{\mathsf{top}}$ and $j\geq 1$ such that $\max_{x}\overline{\operatorname{ht}}_{x}(I)\leq n-j$. Let $G_{I_{\downarrow}}=(V,E\cup\tilde{E})$ denote the induced subgraph of $\Lambda^{\prime}_{n,n}$ on $V=\widehat{\mathcal{V}}_{\mathsf{top}}^{c}(I)$, augmented via edges $\tilde{E}$ connecting every $v\in\mathcal{L}_{-1/2}\cap\Lambda^{\prime}_{n,n}$ to an (arbitrarily chosen) external boundary vertex $v_{*}\in\mathcal{L}_{<0}\setminus\Lambda_{n,n}^{\prime}$. Let $\tilde{\pi}_{I_{\downarrow}}^{(\theta)}$ be the FK model on $G_{I_{\downarrow}}$ with probabilities $p=1-e^{-\beta}$ in $E$ and $\theta$ in $\tilde{E}$, and boundary conditions wired on $\mathcal{L}_{<0}\cap\partial\Lambda^{\prime}_{n,n}$ (and free elsewhere). Then $\Xi_{j}^{\mathsf{fl}}(I)$ and $\Xi_{j}^{\mathsf{dob}}(I)$ from (2.18) satisfy
@@
Our proof crucially uses the fact that the event $\mathcal{I}_{\mathsf{top}}=I$ is measurable w.r.t. the configuration “above $I$” (on $G_{I_{\uparrow}}$ in the notation from the proofs above), whereas it is only “below $I$” (on $G_{I_{\downarrow}}$) where there is a difference between the hard floor and soft floor measures, as $I\subset\mathcal{L}_{\geq 0}$ in both. Thus, the nontrivial portion (“above $I$”) cancels in the ratio $\Xi_{j}^{\mathsf{fl}}/\Xi_{j}^{\mathsf{dob}}$, and leaves us with a ratio of probabilities in two standard FK models. At this point we may use FKG as $\mathcal{I}_{\mathsf{top}}$ reveals a free boundary below it, c.f. $\mathcal{I}_{\mathsf{full}}$ or $\mathcal{I}_{\mathsf{bot}}$ which would reveal a wired boundary below it, and therefore not go in the right direction. (Indeed, the result cannot hold for $\mathcal{I}_{\mathsf{bot}}$, as it would lead to a lower bound of $(1+\varepsilon_{\beta})h_{n}^{*}$ in Theorem 1.1, violating the upper bound $h_{n}^{*}$ given there.)

### v2

@@ 2.1. Random-cluster, FK–Potts and fuzzy Potts models
    Let $\mathcal{V}_{\mathsf{top}}$ denote the vertices of the open cluster of $\omega$ incident to the $\mathsf{red}$ boundary vertices; augment it to $\widehat{\mathcal{V}}_{\mathsf{top}}$ by adding all connected components of $(\mathbb{Z}+\frac{1}{2})^{3}\setminus\mathcal{V}_{\mathsf{top}}$ whose outer vertex boundary is a subset of $\mathcal{V}_{\mathsf{top}}$. Define $\mathcal{V}_{\mathsf{bot}}$, $\widehat{\mathcal{V}}_{\mathsf{bot}}$ analogously. (See Fig. 2.)
@@
By the FK–Potts coupling, under either set of boundary conditions $(\sigma,\omega)\sim{\mathbf{P}}_{n}^{\mathsf{fl}}$ or $(\sigma,\omega)\sim{\mathbf{P}}_{n}^{\mathsf{h}}$, one has $\widehat{\mathcal{V}}_{\mathsf{top}}\subset\widehat{\mathcal{V}}_{\mathsf{red}}\subset\widehat{\mathcal{V}}_{\mathsf{blue}}^{c}$ and $\widehat{\mathcal{V}}_{\mathsf{bot}}\subset\widehat{\mathcal{V}}_{\mathsf{blue}}(\subset\widehat{\mathcal{V}}_{\mathsf{red}}^{c})$. Then the four interfaces are ordered (see Fig. 3) via the partial order on subsets of the vertices [^2] Associate $\mathcal{I}_{\mathsf{blue}}$ and $\mathcal{I}_{\mathsf{bot}}$ with $\widehat{\mathcal{V}}_{\mathsf{blue}}$ and $\widehat{\mathcal{V}}_{\mathsf{bot}}$, resp., and associate $\mathcal{I}_{\mathsf{red}}$ and $\mathcal{I}_{\mathsf{top}}$ with $\widehat{\mathcal{V}}_{\mathsf{red}}^{c}$ and $\widehat{\mathcal{V}}_{\mathsf{top}}^{c}$, resp. (i.e., in all cases, use the part incident to the bottom face of the boundary conditions) to arrive at this partial order. :
@@
Let $\omega\sim\bar{\pi}_{n}^{\mathsf{fl}}$ or $\omega\sim\bar{\pi}_{n}^{\mathsf{h}}$. The below definition applies to both, with the sets $\mathsf{blue}$ boundary vertices”, and $\mathsf{red}$ boundary vertices” interpreted according to the setting. Consider the set $F$ of plaquettes dual to closed edges of $\omega$. The full interface $\mathcal{I}_{\mathsf{full}}$ is the maximal $1$-connected component of $F$ (plaquettes $f,f^{\prime}$ are $1$-adjacent if they share an edge) incident to the boundary plaquettes (plaquettes dual to edges between the $\mathsf{blue}$ and $\mathsf{red}$ boundary vertices).
@@
The above result is stated in [14, Lem. 9],[15, Lem 7.118] for the limit as $m\to\infty$, but its analog was first established for $\bar{\pi}^{\mathsf{h}}_{\Lambda^{\prime}_{n,m}}(\mathcal{I}_{\mathsf{full}}=I)$ in [14, Lem. 7],[15, Lem 7.104] for all $n,m\geq 1$. The proof further applies to any box $\Lambda=([-\frac{n}{2},\frac{n}{2}]^{2}\times[m_{1},m_{2}])\cap(\mathbb{Z}+\frac{1}{2})^{3}$ for $m_{1}\leq 0$ and $m_{2}\geq 1$, with the caveat that the bound on $\mathbf{g}$} in terms of $r(f,I;f^{\prime},I^{\prime})$ (Eq. 2.9) might render Eq. 2.7 useless if $m_{1}$ or $m_{2}$ were taken too close to 0 (as $r$ will be capped at the distance to the boundary).
@@
Let $q\geq 1$, $n\geq 1$ and $m\geq n/2$. There exist $p_{0}<1$ such that, for every $p>p_{0}$ there is some $a=a_{p}>0$ so the following holds,
@@
The above result was stated for the limit $m\to\infty$, and as explained in Remark 2.4, is further applicable to $\Lambda^{\prime}_{n,m}$ with some modifications (the potential pinning effects of the hard floor and ceiling are straightforwardly shown to be negligible when they are positioned at height $\pm cn$; we give a detailed proof in Appendix B for completeness). In our context of $p=1-e^{-\beta}$, and for $\beta$ large enough the result above holds true for $a=\beta-C$ where $C>0$ is an absolute constant.
An important ingredient in the proofs is comparison between the hard floor setting of $\mu_{n}^{\mathsf{fl}}$ and the “soft floor” setting of the conditional distribution of $\mu_{n}^{\mathsf{h}}$ given that $\mathcal{I}_{\mathsf{blue}}$ lies above height 0, namely
@@ 2.2. A logarithmic upper bound on the top interface
As explained by step (2) in the sketch at the beginning of this section, the lower bound argument requires as an input, a logarithmic upper bound on the maximal height of $\mathcal{I}_{\mathsf{top}}$ in the measure $\bar{\pi}_{n}^{\mathsf{fl}}$. Recall that $\mathcal{L}_{\geq h}$ is the slab $\mathbb{R}^{2}\times[h,\infty)$.
@@
Consequently, by Eq. 2.6, the same holds for $\mathcal{I}_{\mathsf{bot}}$ under $\bar{\pi}_{n}^{\mathsf{fl}}$, as well for $\mathcal{I}_{\mathsf{blue}},\mathcal{I}_{\mathsf{red}}$ under $\mu_{n}^{\mathsf{fl}}$.
While the logarithmic upper bound on $\mathcal{I}_{\mathsf{top}}$ implies one for $\mathcal{I}_{\mathsf{blue}}$, our argument for establishing this bound for $\mathcal{I}_{\mathsf{top}}$ goes through first proving such a result directly on $\mathcal{I}_{\mathsf{blue}}$ (Corollary 2.12 below) and then using the rigidity of the $4$ interfaces (Corollary 2.15 below) to conclude the theorem above.
@@ 2.2.1. Upper bound on the blue interface with a soft floor
By Theorems 2.5 and 2.6, a union bound over $x=(x_{1},x_{2})\in[-\frac{n}{2},\frac{n}{2}]^{2}\cap(\mathbb{Z}+\frac{1}{2})^{2}$, as well as on the two events $\{\overline{\operatorname{ht}}_{x}(\mathcal{I}_{\mathsf{full}})\geq h_{0}\}$ and $\{\underline{\operatorname{ht}}_{x}(\mathcal{I}_{\mathsf{full}})\leq-h_{0}\}$ in Theorem 2.5 (which, in our context here of $\bar{\pi}_{n}^{h_{0}}$, address the intersection of $\mathcal{I}_{\mathsf{full}}$ with $\mathcal{L}_{2h_{0}}$ and $\mathcal{L}_{0}$, respectively), gives
@@ 2.2.2. Comparing the soft and hard floor measures
Determining whether $\mathcal{I}_{\mathsf{blue}}\subset\mathcal{L}_{\geq 0}$ must be done with care; a good example to keep in mind in this context is $\sigma$ that colors every $v\in\mathcal{L}_{<0}$ in $\mathsf{blue}$ and every $v\in\mathcal{L}_{>0}$ in $\mathsf{red}$ except for $Q=\{\pm\frac{1}{2}\}^{3}$, where we swap the assignment, coloring $v\in Q\cap\mathcal{L}_{<0}$ in $\mathsf{red}$ and $v\in S\cap\mathcal{L}_{>0}$ in $\mathsf{blue}$. In this example, $0,1]^{2}\times\{-1\}\subset\mathcal{I}_{\mathsf{blue}}$ (in particular $\mathcal{I}_{\mathsf{blue}}\nsubseteq\mathcal{L}_{\geq 0}$) and similarly $0,1]^{2}\times\{1\}\subset\mathcal{I}_{\mathsf{red}}$, since $Q\subset\widehat{\mathcal{V}}_{\mathsf{blue}}^{c}\cap\widehat{\mathcal{V}}_{\mathsf{red}}^{c}$. On the other hand, if we were to also set the color of $(\frac{3}{2},\frac{1}{2},\frac{1}{2})$ to $\mathsf{blue}$, it would connect $Q\cap\mathcal{L}_{>0}$ to $\mathcal{V}_{\mathsf{blue}}$, thus leading to $Q\subset\widehat{\mathcal{V}}_{\mathsf{blue}}$ and to $\mathcal{I}_{\mathsf{blue}}=\mathcal{I}_{\mathsf{red}}\subset\mathcal{L}_{\geq 0}$.
$V_{0}$$V_{1}$$\Gamma$
@@
arriving at $\phi_{n}^{\mathsf{fl}}$, a fuzzy Potts measure on $\Lambda_{n}$ with a hard $\mathsf{blue}$ floor at height 0, a single vertex $v_{\mathsf{red}}$ replacing all the $\mathsf{red}$ boundary vertices in $\mathcal{L}_{>\mathsf{h}}$, and that vertex connected to every formerly $\mathsf{blue}$ boundary condition vertex in $\partial\Lambda_{n,n}\cap\mathcal{L}_{>0}\cap\mathcal{L}_{<\mathsf{h}}$ via edges with interaction strength $J$):
@@ 2.2.3. Rigidity of the four interfaces
Let $\omega\sim\bar{\pi}_{n}^{\mathsf{fl}}$, consider $v\in\Lambda_{n,n}$. Let $\mathcal{C}_{v}$ be the (potentially empty) $(\mathbb{Z}+\frac{1}{2})^{3}$-connected component of the vertex $v$ in $\Lambda_{n,n}\setminus(\widehat{\mathcal{V}}_{\mathsf{top}}\cup\widehat{\mathcal{V}}_{\mathsf{bot}})$. Let $\Gamma_{v}$ be its plaquette boundary, that is, $f\in\Gamma_{v}$ if it is dual to the $(\mathbb{Z}+\frac{1}{2})^{3}$-bond $(u,w)$ with $u\in\mathcal{C}_{v}$ and $w\notin\mathcal{C}_{v}$. Then
@@
Similarly, we can easily bound the differences between the $\mathsf{top}$ and $\mathsf{full}$ interfaces.
Lemma 2.14.
Fix an edge $e\in\Lambda_{n}$ and consider the dual face $f=f_{e}$. Let $C_{f}$ be the (potentially empty) 1-connected component of $f$ in $\mathcal{I}_{\mathsf{full}}\setminus\mathcal{I}_{\mathsf{top}}$. Then,
$\bar{\pi}_{n}^{\mathsf{fl}}(|C_{f}|\geq r)\leq e^{-(\beta-C)r}\,.$
The same statement holds if we define $C_{f}$ to be the 1-connected component of $f$ in $\mathcal{I}_{\mathsf{full}}\setminus\mathcal{I}_{\mathsf{bot}}$.
Proof.
Take the map $\Phi$ that opens every edge dual to $C_{f}$. Since $C_{f}\subset\mathcal{I}_{\mathsf{full}}\setminus\mathcal{I}_{\mathsf{top}}$, this operation keeps the resulting configuration in ${\mathfrak{D}}_{n,n}$. The same short computation as in the proof of Lemma 2.13 immediately implies the desired bound. ∎
@@
Fix $\beta$ large enough. There exists a constant $C>0$ such that for any $k\geq 0$,
$\bar{\pi}_{n}^{\mathsf{fl}}\left(\max_{x}\overline{\operatorname{ht}}_{x}(\mathcal{I}_{\mathsf{full}})>\max_{x}\overline{\operatorname{ht}}_{x}(\mathcal{I}_{\mathsf{bot}})+k\right)\leq Cn^{3}e^{-(\beta-C)k/3}\,$,
and
$\bar{\pi}_{n}^{\mathsf{fl}}\left(\min_{x}\underline{\operatorname{ht}}_{x}(\mathcal{I}_{\mathsf{full}})<\min_{x}\underline{\operatorname{ht}}_{x}(\mathcal{I}_{\mathsf{top}})-k\right)\leq Cn^{3}e^{-(\beta-C)k/3}\,.$
By the ordering of the interfaces (2.6), the distance between the maximum of any of the interfaces $\mathsf{full},\mathsf{top},\mathsf{red},\mathsf{blue},\mathsf{bot}$ satisfies the same bound, and likewise for the minimum.
@@
Every $f\in\mathcal{I}_{\mathsf{top}}\setminus\mathcal{I}_{\mathsf{bot}}$ belongs to the boundary $\Gamma_{v}$ of some component $\mathcal{C}_{v}$ as per the definition in Lemma 2.13. Note that said boundary must also contain some plaquette $f^{\prime}\in\mathcal{I}_{\mathsf{bot}}\setminus\mathcal{I}_{\mathsf{top}}$ (some external boundary vertex of $\mathcal{C}_{v}$ must be in $\widehat{\mathcal{V}}_{\mathsf{bot}}$, or else $\mathcal{C}_{v}$ would be part of $\widehat{\mathcal{V}}_{\mathsf{top}}$), whence $|\Gamma_{v}|$ is an upper bound on the length of a path of plaquettes connecting $f,f^{\prime}$. Hence, the difference between $\max_{x}\overline{\operatorname{ht}}_{x}(\mathcal{I}_{\mathsf{top}})$ and $\max_{x}\overline{\operatorname{ht}}_{x}(\mathcal{I}_{\mathsf{bot}})$ is bounded by $\max_{v}|\Gamma_{v}|$. Similarly, the difference between $\max_{x}\overline{\operatorname{ht}}_{x}(\mathcal{I}_{\mathsf{full}})$ and $\max_{x}\overline{\operatorname{ht}}_{x}(\mathcal{I}_{\mathsf{top}})$ is bounded by $\max_{f}|C_{f}|$. We conclude by Lemmas 2.132.14 and a union bound over the $O(n^{3})$ vertices in $\Lambda_{n}$. ∎
@@
Let $(\sigma,\omega)\sim{\mathbf{P}}_{n}^{\mathsf{fl}}$. By taking $k=(12/\beta)\log n$ in Corollary 2.15, combined with the ordering of the interfaces in Eq. 2.6, with probability at least $1-n^{-1+\varepsilon_{\beta}}$ we have that for every $x$,
@@
[…]
@@
[…]
@@
Our proof crucially uses the fact that the event $\mathcal{I}_{\mathsf{top}}=I$ is measurable w.r.t. the configuration “above II” (on $G_{I_{\uparrow}}$ in the notation from the proofs above), whereas it is only “below II” (on $G_{I_{\downarrow}}$) where there is a difference between the hard floor and soft floor measures, as $I\subset\mathcal{L}_{\geq 0}$ in both. Thus, the nontrivial portion (“above II”) cancels in the ratio $\Xi_{j}^{\mathsf{fl}}/\Xi_{j}^{\mathsf{dob}}$, and leaves us with a ratio of probabilities in two standard FK models. At this point we may use FKG as $\mathcal{I}_{\mathsf{top}}$ reveals a free boundary below it, c.f. $\mathcal{I}_{\mathsf{full}}$ or $\mathcal{I}_{\mathsf{bot}}$ which would reveal a wired boundary below it, and therefore not go in the right direction. (Indeed, the result cannot hold for $\mathcal{I}_{\mathsf{bot}}$, as it would lead to a lower bound of $(1+\varepsilon_{\beta})h_{n}^{*}$ in Theorem 1.1, violating the upper bound $h_{n}^{*}$ given there.)